### v1

# Dyonic RN-like and Taub-NUT-like black holes in Einstein-bumblebee gravity

Abstract
Einstein-bumblebee gravity is one of the simplest vector-tensor theories that realizes spontaneous Lorentz symmetry breaking. In this work, we first construct an exact dyonic Reissner-Nordström-like black hole solution in four dimensions, carrying both electric and magnetic charges and admitting general topological horizons. We then study its thermodynamic properties, and employ the Wald formalism to compute the conserved mass and entropy, thereby establishing the first law of black hole thermodynamics. Furthermore, we generalize these results to Taub-Newman-Unti-Tamburino case and higher dimensions case.
I Introduction
Lorentz symmetry, or Lorentz invariance, is a fundamental postulate of both Standard Model and general relativity (GR). Over the past decades, its validity has been extensively investigated from both theoretical and experimental perspectives [1, 2, 3, 4, 5, 6, 7]. From the theoretical side, it has been suggested that Lorentz invariance might not be an exact symmetry at all energy scales. In this context, a variety of quantum gravity models accommodating Lorentz violation have been proposed, including string theory [8, 9], warped brane worlds [10], and loop quantum gravity [11]. If Lorentz invariance is broken, one naturally expects significant violation at the Planck scale, around $10^{19}$ GeV, while small residual effects could appear at low energies. Furthermore, even if a quantum gravity theory preserves Lorentz invariance at the Planck scale, it may contain tensor fields that acquire nonzero vacuum expectation values (VEVs) at lower energies, thereby spontaneously breaking the symmetry [8]. With steady advances in experimental techniques, the search for low-energy signatures of Lorentz violation remains an active and enduring line of research.
Based on the idea of spontaneous Lorentz symmetry breaking in string theory [8], Colladay and Kostelecký [12, 13, 14] proposed a model-independent framework, known as the Standard Model extension, to systematically incorporate almost all possible Lorentz-violating effects in both the Standard Model and GR. In the gravitational sector, many important modified gravity theories, particularly those involving coupled vector fields, can be formulated within the framework of Standard Model extension. For a comprehensive review, we refer the reader to Ref. [2] and references therein.
Among these Lorentz-violating gravity theories, Einstein-bumblebee gravity [14, 15] represents one of the simplest vector-tensor models that realizes spontaneous Lorentz symmetry breaking. Since its proposal, it has attracted sustained research interest. In particular, Casana et al. [16] constructed an elegant Schwarzschild-like exact black hole solution, which allows one to conveniently study the implications of spontaneous Lorentz symmetry breaking in the gravitational sector through black hole physics and astrophysical phenomena, thereby further stimulating attention to the theory. Subsequently, several exact black hole solutions, analogous to their GR counterparts, have been constructed within Einstein-bumblebee gravity, including Schwarzschild-like black hole with a cosmological constant [17], electrically charged Reissner-Nordström (RN)-like black holes [18], Taub-Newman-Unti-Tamburino (Taub-NUT)-like black holes [19], high dimensional Schwarzschild-like black holes with a cosmological constant [20], as well as a variety of other analytic and numerical solutions [21, 22, 23, 24, 25, 26]. The properties of these black holes have also been investigated [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 59, 60, 61, 62, 58]. Moreover, Einstein-bumblebee gravity has been studied in a variety of other contexts, including compact stars, cosmology, gravitational waves, and generalized formulations, and so on [63, 64, 65, 66, 73, 67, 68, 69, 70, 71, 72, 74, 78, 79, 75, 76, 77, 80, 82, 80, 83, 84, 81, 85, 86]. For a more comprehensive overview of developments in this gravity model, we refer the reader to Ref. [14, 15] and references therein.
In this work, we construct exact solutions for a class of dyonic RN-like black holes in four dimensions, carrying both electric and magnetic charges and admitting general topological horizons, and study their thermodynamic properties within Einstein-bumblebee gravity. We further extend the study to include the Taub-NUT case and to higher-dimensional spacetimes. The motivations are as follows. First, both electrically and magnetically charged black holes have long been of strong theoretical and astrophysical interest [87, 88, 89, 90, 91, 92, 93], constructing exact dyonic RN-like solutions in Einstein-bumblebee gravity provides a convenient setting for exploring the implications of spontaneous Lorentz symmetry breaking. Second, the thermodynamics of neutral black holes in this theory exhibits notable subtleties: the usual definitions of mass (Komar and Arnowitt-Deser-Misner mass) and entropy (Wald entropy) fail to yield a consistent first law, as reported in the literatures, such as [49, 19]. Consistency is restored only by employing more general methods, such as the Wald formalism [94, 95], to properly define the conserved mass charge and entropy. Extending such an analysis to cases with Maxwell fields, topological horizons, and higher dimensions is therefore both necessary and well motivated. Third, although previous works have reported four-dimensional electrically charged RN-like black hole solutions with spherical horizons [18], which have been widely applied in black hole physics and astrophysics [18, 59, 60, 61, 62], the previous Einstein-bumblebee-Maxwell (EbM) theory does not admit a purely magnetic black hole, while the magnetic charge of the dyonic black holes is not an independent integration constant, but fixed by the coupling constant of the theory. In contrast, we employ an extended framework within the EbM theory family [14, 15, 80] that admits not only purely electric or magnetic black holes, but a well-defined exact dyonic black hole solution, where the charges are true integration constants, independent of the coupling constants of the theory. This provides a foundation for further investigation of the physical properties and potential observational signatures of dyonic black holes, as well as for exploring more complex black hole solutions within a consistent Einstein-bumblebee gravity framework. In particular, the construction of exact dyonic Taub-NUT-like black hole solutions and higher-dimensional dyonic RN-like solutions within this setup further illustrates the internal consistency and robustness of the specific EbM model.
This paper is organized as follows: In Sec. II, we briefly review the EbM theory family and its general equations of motion (EOMs). In Sec. III, we derive the four-dimensional dyonic RN-like topological black holes in a particular theory. In Sec. IV, we study the thermodynamics of these black holes and employ the Wald formalism to compute the conserved mass and entropy, thereby establishing the first law of black hole thermodynamics. Furthermore, in Sec. V, we construct the dyonic Taub-NUT-like solution and examine its thermodynamic properties using the Wald formalism. In Sec. VI we generalize previous RN-like results to higher dimensions. Finally, We presents a summary and discussion of the work in Sec. VII.
II The theory and its EOMs
In this section, we briefly review the EbM theory. The total action $\cal S$ in $D$-dimensions ($D\geq 4$) can be expressed as
${\cal S}=\frac{1}{2\kappa}\int d^{D}x\sqrt{-g}\left(L_{1}+L_{2}\right)\,$, (1)
with $\kappa=8\pi G/c^{4}$, where $G$ and $c$ denote the gravitational constant and velocity of light. The Lagrangian densities $L_{1}$ for Einstein-bumblebee gravity sector is given by [14, 15]
$L_{1}=R+\gamma R_{\mu\nu}B^{\mu}B^{\nu}-2\kappa\left(\frac{1}{4}B_{\mu\nu}B^{\mu\nu}+V(B_{\mu}B^{\mu})\right)\,$, (2)
where $g$ denotes the determinant of the metric $g_{\mu\nu}$, $R$ the Ricci scalar, and $R_{\mu\nu}$ the Ricci tensor. As a class of vector-tensor model, Einstein–bumblebee gravity introduces a vector field $B_{\mu}$, commonly referred to as the bumblebee field, which triggers spontaneous Lorentz symmetry breaking. Its field strength $B_{\mu\nu}$ is defined as
$B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}\,.$ (3)
The parameter $\gamma$ denotes the coupling constant associated with the non-minimal interaction between the bumblebee field and the Ricci tensor. The potential $V$ should be minimized to obtain a stable vacuum of spacetime when the bumblebee field acquires a nonzero VEV $b_{\mu}$, i.e., $\langle B_{\mu}\rangle=b_{\mu}$. This nonzero VEV introduces a preferred direction in spacetime, thereby leading to the spontaneous breaking of Lorentz symmetry. Consequently, the potential can generally be expressed as
$V\equiv V(B_{\mu}B^{\mu}\pm b^{2})\,$, (4)
where $b^{2}$ represents a real positive constant. Following discussions of symmetry-breaking potentials in the literature [14, 15, 16], the minimum of $V$ is typically chosen to vanish, which implies
$V(B_{\mu}B^{\mu}\pm b^{2})\big|_{B_{\mu}=b_{\mu}}$ $=$ $0\,$, (5)
$V^{\prime}(B_{\mu}B^{\mu}\pm b^{2})\big|_{B_{\mu}=b_{\mu}}$ $=$ $0\,$, (6)
where $V^{\prime}(x)\equiv dV(x)/dx$ and the $b_{\mu}$ is determined by
$b_{\mu}b^{\mu}\pm b^{2}=0\,.$ (7)
The $\pm$ signs in Eq. (7) determine whether the field $b_{\mu}$ is timelike or spacelike. The electromagnetic sector $L_{2}$ is governed by Maxwell theory, with the Maxwell field also exhibiting a non-minimal coupling to the bumblebee field [18, 73, 80, 96, 97], which is given by
$L_{2}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\gamma_{1}B_{\lambda}B^{\lambda}F_{\mu\nu}F^{\mu\nu}+\gamma_{2}B^{\mu}F_{\mu\nu}B_{\lambda}F^{\lambda\nu}\,$, (8)
where the Maxwell field is denoted by $A_{\mu}$, with its field strength defined as
$F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}\,.$ (9)
Here, $\gamma_{1}$ and $\gamma_{2}$ are coupling constants characterizing the non-minimal interactions between the Maxwell field and the bumblebee field. The EOMs can be derived by varying the action with respect to the gravitational field $g_{\mu\nu}$, the bumblebee field $B_{\nu}$, and the Maxwell field $A_{\nu}$. The resulting EOMs, denoted by $E_{\mu\nu}$, $E_{B}^{\nu}$ and $E_{A}^{\nu}$, are given by
$E_{\mu\nu}$ $\equiv$ $G_{\mu\nu}+\gamma\bigg[-\frac{1}{2}g_{\mu\nu}R_{\rho\sigma}B^{\rho}B^{\sigma}+2R^{\rho}{}_{(\mu}B_{\nu)}B_{\rho}+\frac{1}{2}g_{\mu\nu}\nabla_{\rho}\nabla_{\sigma}(B^{\rho}B^{\sigma})$ (10)
$+\frac{1}{2}\Box(B_{\mu}B_{\nu})-\nabla_{\rho}\nabla_{(\mu}(B_{\nu)}B^{\rho})\bigg]-\frac{\kappa}{2}\left[2B_{\mu\lambda}B_{\nu}{}^{\lambda}-\frac{1}{2}g_{\mu\nu}B_{\rho\sigma}B^{\rho\sigma}\right]$
$+2\kappa\left[\frac{1}{2}g_{\mu\nu}V-\frac{\partial V}{\partial X}B_{\mu}B_{\nu}\right]-\frac{1}{4}\left[2F_{\mu\lambda}F_{\nu}{}^{\lambda}-\frac{1}{2}g_{\mu\nu}F_{\rho\sigma}F^{\rho\sigma}\right]$
$+\gamma_{1}\left[B_{\rho}B^{\rho}\left(2F_{\mu\lambda}F_{\nu}{}^{\lambda}-\frac{1}{2}g_{\mu\nu}F_{\rho\sigma}F^{\rho\sigma}\right)+B_{\mu}B_{\nu}F_{\rho\sigma}F^{\rho\sigma}\right]$
$+\gamma_{2}\left[-\frac{1}{2}g_{\mu\nu}B^{\rho}B_{\lambda}F_{\rho\sigma}F^{\lambda\sigma}+2B_{\sigma}B_{(\mu}F_{\nu)\rho}F^{\sigma\rho}+B^{\rho}B^{\sigma}F_{\mu\rho}F_{\nu\sigma}\right]=0\,$,
$E_{B}^{\nu}$ $\equiv$ $\gamma R^{\mu\nu}B_{\mu}+\kappa\nabla_{\mu}B^{\mu\nu}-2\kappa\frac{\partial V}{\partial X}B^{\nu}+\gamma_{1}B^{\nu}F_{\rho\sigma}F^{\rho\sigma}+\gamma_{2}B^{\mu}F_{\mu\lambda}F^{\nu\lambda}=0\,$, (11)
$E_{A}^{\nu}$ $\equiv$ $\nabla_{\mu}\left[(1-4\gamma_{1}B_{\lambda}B^{\lambda})F^{\mu\nu}+4\gamma_{2}B_{\lambda}F^{\lambda[\mu}B^{\nu]}\right]=0\,$, (12)
where $G_{\mu\nu}\equiv R_{\mu\nu}-Rg_{\mu\nu}/2$ is the Einstein tensor, $X=B_{\mu}B^{\mu}\pm b^{2}$, $\nabla_{\mu}$ is the covariant derivative, $\Box=\nabla_{\mu}\nabla^{\mu}$ is the d’Alembert operator. Parentheses $(\mu\nu)$ and square brackets $\mu\nu]$ indicate symmetrization and antisymmetrization over the enclosed indices, respectively.
III Dyonic RN-like topological black holes in four dimensions
In this section, we construct the dyonic solution of (1). A comparison of the coupling constants in the Maxwell sector of (8) shows that the bumblebee theory considered here differs from that studied in [18]. Moreover, we will illustrate the advantages of the theory developed in this work from the perspective of its dyonic solutions.
III.1 Exact dyonic solution
We now construct the static dyonic RN-like black hole solution with general topological horizons in $D=4$ dimensions. The most general static ansatz for the metric $g_{\mu\nu}$ and the Maxwell field potential $A_{\mu}$ can be written as
$ds^{2}$ $=$ $-h(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega_{2,k}^{2}\,$, (13)
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\phi(r)dt+pud\varphi\,$, (14)
where $d\Omega_{2,k}^{2}={du^{2}}/{(1-ku^{2})}+(1-ku^{2})d\varphi^{2}$ with k=1,0,1k=1,0,-1, corresponding to the metric of the unit 2-sphere, the 2-torus or the unit hyperbolic 2-space, respectively. The constant $p$ denotes the magnetic charge parameter. For the bumblebee field $B_{\mu}$, following Ref. [16], we consider a nonzero VEV $b_{\mu}$ oriented along a radial direction, taking the form
$B_{{{\scriptscriptstyle(1)}}}=b_{r}(r)dr\,.$ (15)
Imposing the condition $b_{\mu}b^{\mu}=b^{2}=\textup{const.}$ leads to
$b_{r}=\frac{b}{\sqrt{f}}\,$, (16)
which ensures that the bumblebee field strength vanishes, $b_{\mu\nu}=0$, and that the potential, which may be chosen as $V=(B_{\mu}B^{\mu}-b^{2})^{2}$, satisfies the vacuum conditions in Eqs. (5)–(6). Next, substituting Eqs. (13)–(16) into EOM (12), the $t$-component of $E_{A}^{\mu}$ yields
$\left(\frac{r^{2}\sqrt{h}}{\sqrt{f}}\ \frac{(1-2b^{2}(2\gamma_{1}+\gamma_{2}))f\phi^{\prime}}{4h}\right)^{\prime}=0\,$, (17)
where the prime denotes differentiation with respect to $r$. This equation can be integrated once to give
$\phi^{\prime}=\frac{q\sqrt{h}}{r^{2}\sqrt{f}}\,$, (18)
where $q$ is an integration constant associated with the electric charge. Then, substituting Eqs. (13)–(16) and (18) into EOMs (10)–(11), we obtain the nonzero components
$E^{t}{}_{t}$ $\equiv$ $\frac{1+b^{2}\gamma}{r}\left[f^{\prime}+\frac{f}{r}+\frac{p^{2}\left(1-4b^{2}\gamma_{1}\right)+q^{2}\left(1-2b^{2}(2\gamma_{1}+\gamma_{2})\right)-4kr^{2}}{4r^{3}\left(b^{2}\gamma+1\right)}\right]=0\,$, (19)
$E^{r}{}_{r}$ $\equiv$ $-\frac{b^{2}\gamma f}{2h}\Bigg[h^{\prime\prime}-\frac{h^{\prime 2}}{2h}+h^{\prime}\left(\frac{f^{\prime}}{2f}-\frac{2\left(b^{2}\gamma+1\right)}{b^{2}\gamma r}\right)+\frac{2hf^{\prime}}{rf}-\frac{2h\left(b^{2}\gamma+1\right)}{b^{2}\gamma r^{2}}$ (20)
$+\frac{h\left(q^{2}\left(6b^{2}(2\gamma_{1}+\gamma_{2})-1\right)-p^{2}\left(1+4b^{2}\gamma_{1}\right)+4kr^{2}\right)}{2b^{2}\gamma r^{4}f}\Bigg]=0\,$,
$E^{u}{}_{u}$ $=$ $E^{\varphi}{}_{\varphi}\equiv\frac{\left(b^{2}\gamma+1\right)f}{2h}\Bigg[h^{\prime\prime}-\frac{h^{\prime 2}}{2h}+\left(\frac{f^{\prime}}{2f}+\frac{1}{r}\right)h^{\prime}+\frac{hf^{\prime}}{rf}$ (21)
$+\frac{h\left(q^{2}\left(2b^{2}(2\gamma_{1}+\gamma_{2})-1\right)-p^{2}\left(1-4b^{2}\gamma_{1}\right)\right)}{2r^{4}\left(b^{2}\gamma+1\right)f}\Bigg]=0\,$,
$E^{r}_{B}$ $\equiv$ $-\frac{b\gamma f^{\frac{3}{2}}}{h}\left[h^{\prime\prime}-\frac{h^{\prime 2}}{2h}+\frac{f^{\prime}h^{\prime}}{2f}+\frac{2hf^{\prime}}{rf}-\frac{2h\left(2\gamma_{1}p^{2}-q^{2}(2\gamma_{1}+\gamma_{2})\right)}{\gamma r^{4}f}\right]=0\,.$ (22)
From Eq. (19), one can directly integrate to obtain
$f=\frac{1}{1+b^{2}\gamma}\left[k-\frac{m}{r}+\frac{(1-2b^{2}(2\gamma_{1}+\gamma_{2}))q^{2}}{4r^{2}}+\frac{\left(1-4b^{2}\gamma_{1}\right)p^{2}}{4r^{2}}\right]\,$, (23)
where $m$ is an integration constant related to the mass. Solving Eq. (20) for $h^{\prime\prime}$, and then substituting it together with Eq. (23) into Eq. (22), we obtain
$h=k-\frac{m}{r}+\frac{(1-2b^{2}(2\gamma_{1}+\gamma_{2}))q^{2}}{4r^{2}}+\frac{\left(1-4b^{2}\gamma_{1}\right)p^{2}}{4r^{2}}\,.$ (24)
Then, substituting Eqs. (23)–(24) back into Eq. (20), the existence of nonzero electric and magnetic charges $q$ and $p$ requires
$\gamma_{1}=\frac{\gamma}{4(2+3b^{2}\gamma)}\,,\quad\gamma_{2}=-\frac{2\gamma(1+b^{2}\gamma)}{(2+b^{2}\gamma)(2+3b^{2}\gamma)}\,.$ (25)
Finally, inserting Eqs. (23)–(25) into Eq. (21), one finds that the equation is automatically satisfied. Combining with Eq. (18) and rescaling $q\to q/\sqrt{1+b^{2}\gamma}$, the final solution reads
$h$ $=$ $k-\frac{m}{r}+\frac{q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{\left(1+\ell\right)p^{2}}{2\left(2+3\ell\right)r^{2}}\,,\quad f=\frac{h}{1+\ell}\,,\quad\phi=\frac{q}{r}\,$, (26)
where the Lorentz violating parameter $\ell=b^{2}\gamma$, and the couplings constants are
$\gamma_{1}=\frac{\gamma}{4(2+3\ell)}\,,\quad\gamma_{2}=-\frac{2\gamma(1+\ell)}{(2+\ell)(2+3\ell)}\,.$ (27)
Note that we have chosen the potential $\phi$ to vanish at infinity. In the limit $\ell\to 0$, the solution reduces to the standard dyonic RN black hole with a generic topological horizon. Alternatively, in the absence of both electric and magnetic charges with $k=1$, the solution reduces to a Schwarzschild-like black hole, which has been extensively discussed in Ref. [16]. The Kretschmann scalar, i.e., the square of the Riemann tensor, can be easily calculated as
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ $=$ $\frac{1}{r^{8}}\left[\frac{14p^{4}}{(3\ell+2)^{2}}+\frac{28p^{2}q^{2}}{(\ell+1)(\ell+2)(3\ell+2)}+\frac{14q^{4}}{(\ell+1)^{2}(\ell+2)^{2}}\right]$ (28)
$-\frac{m}{r^{7}}\left[\frac{24p^{2}}{(\ell+1)(3\ell+2)}+\frac{24q^{2}}{(\ell+1)^{2}(\ell+2)}\right]$
$+\frac{1}{r^{6}}\left[-\frac{4kp^{2}\ell}{(\ell+1)(3\ell+2)}-\frac{4kq^{2}\ell}{(\ell+1)^{2}(\ell+2)}+\frac{12m^{2}}{(\ell+1)^{2}}\right]$
$+\frac{8km\ell}{r^{5}(\ell+1)^{2}}+\frac{4k^{2}\ell^{2}}{r^{4}(\ell+1)^{2}}\,$,
which exhibits the similar singularity structure as the dyonic RN case, but clearly deviates from the corresponding Kretschmann invariant when $\ell\neq 0$.
III.2 The difference from [18]
In [18], the authors identified a set of static solutions in the purely electric case that differ from those obtained in this work. After choosing the coupling constants
$\gamma_{1}=-\frac{\gamma}{4(\ell+2)}\,,\qquad\gamma_{2}=0\,$, (29)
the metric $g_{\mu\nu}$, Maxwell field $A_{{{\scriptscriptstyle(1)}}}$, and bumblebee field $B_{{{\scriptscriptstyle(1)}}}$ take the following forms
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}\,,\qquad f=\frac{h}{1+\ell}\,,\qquad\phi=\frac{\sqrt{\ell+1}q}{r}\,,\qquad b_{r}=\frac{b}{\sqrt{f}}\,.$ (30)
Here, we correct a typo in the Maxwell field appearing in [18]. Because the coupling constant $\gamma_{1}$ in this solution differ from that considered in (27), the solution obtained in this work corresponds to a black hole in a distinct bumblebee theory.
Similarly, the above solution can be generalized to the dyonic case
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\phi(r)dt+pud\varphi\,.$ (31)
To obtain an analytical solution, we find that the function $h$ and the coupling constant $\gamma_{2}$ must be adjusted to
$h$ $=$ $k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}}{2(\ell+2)^{2}r^{2}}\,,\qquad\gamma_{2}=-\frac{2\gamma(\ell+1)}{(\ell+2)^{2}}\frac{p^{2}}{q^{2}}\,.$ (32)
We find that in this case, the coupling constant $\gamma_{2}$ depends on the integration constants $p$ and $q$. This indicates that the bumblebee theory (29) considered in [18] cannot be directly extended to arbitrary dyonic black hole solutions, as the coupling constant would otherwise depend on the integration constants. Of course, for the special case $p=q$, such an extension remains possible.
Of course, we can also redefine the parameters $p\to pq$ to make
$\gamma_{2}=-\frac{2\gamma(\ell+1)}{(\ell+2)^{2}}\,$, (33)
independent of the integration constants. However, in this case, the expressions for $A_{{{\scriptscriptstyle(1)}}}$ and $h$
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\phi(r)dt+pqud\varphi\,,\qquad h=k-\frac{m}{r}+\frac{(1+\ell)q^{2}}{2\left(2+\ell\right)r^{2}}+\frac{(\ell+1)(3\ell+2)p^{2}q^{2}}{2(\ell+2)^{2}r^{2}}\,$, (34)
reveal that a purely magnetic black hole cannot be accommodated. Hence, in terms of the feasibility of extending to dyonic solutions, the theory considered here (27) is better suited. In the following Sections V and VI, this approach can be further extended to Taub-NUT case and arbitrary even dimensions, yielding consistent and well-defined dyonic solutions and coupling constants.
IV Thermodynamics
Now we turn to the thermodynamics of the black hole. The event horizon, located at $r=r_{h}$, is determined by the condition $f(r_{h})=0$, where the timelike Killing vector $\xi=\partial_{t}$ becomes null, i.e., $\xi_{\mu}\xi^{\mu}\big|_{r=r_{h}}=0$. It is convenient to express the constant $m$ in terms of $r_{h}$, namely
$m=kr_{h}+\frac{p^{2}(\ell+1)}{2r_{h}(3\ell+2)}+\frac{q^{2}}{2r_{h}(\ell+2)}\,.$ (35)
The black hole temperature $T$ is defined in terms of the surface gravity $K$, which is computed from the Killing vector $\xi$ as
$K^{2}=-\frac{\nabla^{\mu}\xi^{\nu}\nabla_{\mu}\xi_{\nu}}{2}\Big|_{r=r_{h}}\,,\quad T=\frac{K}{2\pi}\,.$ (36)
Accordingly, the temperature reads
$T=\frac{h^{\prime}\left(r_{h}\right)}{4\pi\sqrt{\ell+1}}=\frac{k}{4\pi\sqrt{\ell+1}r_{h}}-\frac{q^{2}}{8\pi\sqrt{\ell+1}(\ell+2)r_{h}^{3}}-\frac{p^{2}\sqrt{\ell+1}}{8\pi(3\ell+2)r_{h}^{3}}\,.$ (37)
Next, in order to obtain the electric charge $Q_{e}$, it is convenient to introduce an antisymmetric tensor ${\cal F}^{\mu\nu}$
${\cal F}^{\mu\nu}=(1-4\gamma_{1}B_{\lambda}B^{\lambda})F^{\mu\nu}+4\gamma_{2}B_{\lambda}F^{\lambda[\mu}B^{\nu]}\,$, (38)
to rewrite the Maxwell field EOM (12) in a compact form:
$E_{A}^{\nu}\equiv\nabla_{\mu}{\cal F}^{\mu\nu}=0\,.$ (39)
Equivalently, the above expression can be cast in the language of differential forms as
$d*\mathcal{F}_{{{\scriptscriptstyle(2)}}}=0\,,\quad*\mathcal{F}_{{{\scriptscriptstyle(2)}}}=\frac{1}{2!2!}\epsilon_{\mu\nu\rho\sigma}\mathcal{F}^{\rho\sigma}dx^{\mu}\wedge dx^{\nu}\,$, (40)
where the subscript $(n)$ denotes an $n$-form, $d$ is the exterior derivative, * the Hodge dual, and the Levi-Civita tensor $\epsilon_{\mu\nu\rho\sigma}$ is defined as $\epsilon_{\mu\nu\rho\sigma}=\sqrt{-g}\varepsilon_{\mu\nu\rho\sigma}$ with $\varepsilon_{0123}=1$. In terms of the Eq. (40), the electric charge $Q_{e}$ is then defined as
$Q_{e}=-\frac{1}{2\kappa}\int_{\Omega_{2,k}}*\mathcal{F}_{{{\scriptscriptstyle(2)}}}=\frac{1}{2\kappa}\int_{\Omega_{2,k}}\sqrt{-g}\varepsilon_{tru\varphi}\mathcal{F}^{rt}dud\varphi=\frac{qw_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}\,$, (41)
where $w_{2}=\int_{\Omega_{2,k}}dud\varphi$. For $k=1$, corresponding to the unit 2-sphere, one has $w_{2}=4\pi$. The electric potential $\Phi_{e}$ is defined through $A_{{{\scriptscriptstyle(1)}}}$
$\Phi_{e}=i_{\xi}A_{{{\scriptscriptstyle(1)}}}|^{r=r_{h}}_{r\rightarrow\infty}=\xi^{\mu}A_{\mu}|^{r=r_{h}}_{r\rightarrow\infty}=\frac{q}{r_{h}}\,$, (42)
where $i_{\xi}$ denotes a contraction of $\xi^{\mu}$ on the index of the $1$-form $A_{{{\scriptscriptstyle(1)}}}$. The Maxwell field strength $F_{{{\scriptscriptstyle(2)}}}=dA_{{{\scriptscriptstyle(1)}}}$ satisfies the Bianchi identity $dF_{{{\scriptscriptstyle(2)}}}=0$, which allows the magnetic charge $Q_{m}$ to be defined in terms of it
$Q_{m}=\frac{1}{2\kappa}\int_{\Omega_{2,k}}F_{{{\scriptscriptstyle(2)}}}=\frac{1}{2\kappa}\int_{\Omega_{2,k}}F_{u\varphi}dud\varphi=\frac{w_{2}p}{2\kappa}\,.$ (43)
According to [98], since $\mathcal{F}_{{{\scriptscriptstyle(2)}}}$ is closed when the on-shell condition is satisfied (i.e. the EOMs are satisfied), we can define a corresponding 1-form $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ satisfying $\mathcal{F}_{{{\scriptscriptstyle(2)}}}=d\mathcal{A}_{{{\scriptscriptstyle(1)}}}$. The $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ is given by
$\mathcal{A}_{{{\scriptscriptstyle(1)}}}=\frac{2p(\ell+1)^{3/2}}{r(3\ell+2)}dt-\frac{2q\sqrt{\ell+1}}{(\ell+2)}ud\varphi\,.$ (44)
Similarly, the magnetic potential $\Phi_{m}$ can therefore be defined in terms of $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$
$\Phi_{m}=i_{\xi}\mathcal{A}_{{{\scriptscriptstyle(1)}}}|^{r=r_{h}}_{r\rightarrow\infty}=\xi^{\mu}\mathcal{A}_{\mu}|^{r=r_{h}}_{r\rightarrow\infty}=\frac{2p(\ell+1)^{3/2}}{(3\ell+2)r_{h}}\,.$ (45)
To complete the thermodynamic quantities, we now turn to the mass and entropy of the black hole. The Komar mass associated with the timelike Killing vector $\xi$ is
$M_{\textup{K}}=-\frac{1}{\kappa}\int_{\Omega_{2,k}}*d\xi\big|_{r\to\infty}=\frac{2}{\kappa}\int_{\Omega_{2,k}}dud\varphi\sqrt{-g}\varepsilon_{tru\varphi}\nabla^{r}\xi^{t}\big|_{r\to\infty}=\frac{mw_{2}}{\kappa\sqrt{1+\ell}}\,.$ (46)
The Wald entropy formula [94, 95] reads
$S_{\textup{W}}=-\frac{\pi}{\kappa}\int_{\Omega_{2,k}}r^{2}\frac{\partial L_{1}}{\partial R_{\mu\nu\rho\sigma}}\epsilon_{\mu\nu}\epsilon_{\rho\sigma}\bigg|_{r\to r_{h}}=\frac{w_{2}\pi}{\kappa}(\ell+2)r_{h}^{2}\,.$ (47)
where the binormal is $\epsilon_{\mu\nu}=\sqrt{h/f}(\delta_{\mu}^{t}\delta_{\nu}^{r}-\delta_{\mu}^{r}\delta_{\nu}^{t})$. However, the Komar mass and Wald entropy, together with the temperature, electric and magnetic charges, and their corresponding potentials, do not satisfy the first law of black hole thermodynamics:
$\delta M_{\textup{K}}\neq T\delta S_{\textup{W}}+\Phi_{e}\delta Q_{e}+\Phi_{m}\delta Q_{m}\,.$ (48)
Similar discrepancies have also been observed in the neutral cases, as discussed in Refs. [49, 19]. In fact, similar violations of the first law due to naive application of definitions have also been reported in other theories, such as four-dimensional gauged supergravities [99, 100, 101], Horndeski gravity [102, 103], and related setups [104, 105]. Motivated by this, we employ the Wald formalism [94, 95] to consistently recompute the mass and entropy.
IV.1 Wald formalism
In 1993, Wald developed the covariant phase space formalism [94, 95]. Within this framework, he demonstrated that in a diffeomorphism-invariant theory of gravity, the Killing vector $\xi^{\mu}$ generates all conserved charges, thereby unifying these quantities into a single differential relation, i.e., the first law of black hole mechanics. In other words, the first law emerges as a natural consequence of Noether’s theorem, ensuring that any diffeomorphism-invariant theory of gravity necessarily admits a self-consistent black hole thermodynamics.
A direct consequence of the Wald formalism is the Wald entropy formula, which extends the Bekenstein-Hawking entropy and possesses remarkable universality. In higher-derivative gravity theories, black hole entropy is no longer proportional to the horizon area. The Wald entropy formula, however, correctly accounts for these corrections [106, 107, 108] and is thus widely used in many black hole–related effective theories [110, 104, 105, 109]. Thus far, the Wald formalism and the Wald entropy formula have been applied to a broad range of extended gravity theories, including tensor-scalar theories [111, 112, 102, 103], Maxwell theory and its extensions [113, 114, 115, 113, 106, 99, 100], as well as the neutral bumblebee theory [49, 19].
In fact, the Wald entropy formula is simply a corollary of the Wald formalism and can be obtained from it only under specific conditions. Indeed, previous studies of Horndeski theory and bumblebee theory have demonstrated that the Wald entropy formula fails to provide a black hole entropy consistent with the first law of thermodynamics [49, 19, 102, 103]. More precisely, the dependence of the Noether charge on the Killing vector involves only $\xi^{\mu}$ and $\nabla^{\mu}\xi^{\nu}$. On the black hole horizon, if all fields are smooth, one has $\xi\to 0$ and $\nabla^{\mu}\xi^{\nu}\to K\epsilon^{\mu\nu}$, from which the Wald entropy formula follows. However, the $B_{\mu}$ field (15, 16) diverges on the black hole horizon, rendering the Wald entropy formula inapplicable. Hence, the complete thermodynamics must be derived directly from the original Wald formalism.
Consider a diffeomorphism invariant theory whose action $\cal S$} in the language of differential form is given by
${\cal S}[\psi]=\frac{1}{2\kappa}\int\mathbf{L}(\psi)\,$, (49)
where $\psi$ collectively denotes all dynamical fields in the system. The Lagrangian $D$-form $\mathbf{L}$ is the Hodge dual of the Lagrangian density $L=L_{1}+L_{2}$, namely,
$\mathbf{L}=*L=\frac{\sqrt{-g}}{D!}\varepsilon_{\alpha_{1}\alpha_{2}\dots\alpha_{D}}Ldx^{\alpha_{1}}\wedge dx^{\alpha_{2}}\wedge\dots\wedge dx^{\alpha_{D}}\,.$ (50)
The variation of the Lagrangian $\mathbf{L}$ takes the general form
$\delta\mathbf{L}=\mathbf{E}_{\psi}\delta\psi+d\mathbf{\Theta}(\psi,\delta\psi)\,$, (51)
where $\mathbf{E}_{\psi}$ represents the EOM, and $\mathbf{\Theta}$ is the symplectic potential (D1)(D-1)-form. Explicitly,
$\mathbf{\Theta}$ $=$ $\frac{\sqrt{-g}}{(D-1)!}\varepsilon_{\mu\alpha_{1}\alpha_{2}\dots\alpha_{D-1}}\Theta^{\mu}dx^{\alpha_{1}}\wedge dx^{\alpha_{2}}\wedge\dots\wedge dx^{\alpha_{D-1}}\,$, (52)
$i_{\xi}\mathbf{\Theta}$ $=$ $\frac{\sqrt{-g}}{(D-2)!}\varepsilon_{\mu\nu\alpha_{1}\alpha_{2}\dots\alpha_{D-2}}\Theta^{\mu}\xi^{\nu}dx^{\alpha_{1}}\wedge dx^{\alpha_{2}}\wedge\dots\wedge dx^{\alpha_{D-2}}\,.$ (53)
From $\mathbf{\Theta}$, one can define the symplectic current (D1)(D-1)-form $\mathbf{\omega}$ as
$\mathbf{\omega}(\psi,\delta_{1}\psi,\delta_{2}\psi)=\delta_{1}\mathbf{\Theta}(\psi,\delta_{2}\psi)-\delta_{2}\mathbf{\Theta}(\psi,\delta_{1}\psi)\,.$ (54)
We next specialize to a variation induced by an infinitesimal diffeomorphism generated by an arbitrary vector field $\xi^{\mu}$, the dynamical fields transform as
$\delta_{\xi}\psi=\mathscr{L}_{\xi}\psi\,$, (55)
with $\mathscr{L}_{\xi}$ representing the Lie derivative along $\xi^{\mu}$. The corresponding variation of the Lagrangian $\mathbf{L}$ can likewise be expressed in two equivalent forms. On the one hand, from the general variation formula we have
$\delta_{\xi}\mathbf{L}=\mathbf{E}_{\psi}\delta_{\xi}\psi+d\mathbf{\Theta}(\psi,\delta_{\xi}\psi)\,$, (56)
while on the other hand, using the diffeomorphism invariance of the theory, it is given by the Lie derivative of $\mathbf{L}$, namely
$\mathscr{L}_{\xi}\mathbf{L}=i_{\xi}d\mathbf{L}+d(i_{\xi}\mathbf{L})=d(i_{\xi}\mathbf{L})\,$, (57)
where we have used $d\mathbf{L}=0$ since the $\mathbf{L}$ is a $D$-form on a $D$-dimensional manifold. One can introduce a Noether current (D1)(D-1)-form $\mathbf{J}_{\xi}$, defined by
$\mathbf{J}_{\xi}=\mathbf{\Theta}(\psi,\delta_{\xi}\psi)-i_{\xi}\mathbf{L}\,.$ (58)
Combining Eqs. (56) and (57) gives
$d\mathbf{J}_{\xi}=-\mathbf{E}_{\psi}\delta_{\xi}\psi\,$, (59)
which vanishes on-shell. Consequently, there exists a Noether charge (D2)(D-2)-form $\mathbf{Q}_{\xi}$ such that
$\mathbf{J}_{\xi}=d\mathbf{Q}_{\xi}\,.$ (60)
The Noether charge can be expressed as
$\mathbf{Q}_{\xi}=\frac{\sqrt{-g}}{(D-2)!2!}\varepsilon_{\mu\nu\alpha_{1}\alpha_{2}\dots\alpha_{D-2}}Q_{\xi}^{\mu\nu}dx^{\alpha_{1}}\wedge dx^{\alpha_{2}}\wedge\dots\wedge dx^{\alpha_{D-2}}\,.$ (61)
Varying the Noether current $\mathbf{J}_{\xi}$, we obtain
$\delta\mathbf{J}_{\xi}$ $=$ $\delta\mathbf{\Theta}(\psi,\delta_{\xi}\psi)-i_{\xi}\delta\mathbf{L}$ (62)
$=$ $\delta\mathbf{\Theta}(\psi,\delta_{\xi}\psi)-i_{\xi}d\mathbf{\Theta}(\psi,\delta\psi)-i_{\xi}(\mathbf{E}_{\psi}\delta\psi)$
$=$ $\delta\mathbf{\Theta}(\psi,\mathscr{L}_{\xi}\psi)-\mathscr{L}_{\xi}\mathbf{\Theta}(\psi,\delta\psi)+d(i_{\xi}\mathbf{\Theta}(\psi,\delta\psi))\,.$
where we have used the on-shell condition, $\mathbf{E}_{\psi}=0$. According to Eq. (54), the first two terms of the right hand side of the Eq. (62) can be identified as the symplectic current (D1)(D-1)-form $\mathbf{\omega}$:
$\mathbf{\omega}(\psi,\delta\psi,\mathscr{L}_{\xi}\psi)$ $=$ $\delta\mathbf{\Theta}(\psi,\mathscr{L}_{\xi}\psi)-\mathscr{L}_{\xi}\mathbf{\Theta}(\psi,\delta\psi)$ (63)
$=$ $\delta\mathbf{J}_{\xi}-d(i_{\xi}\mathbf{\Theta}(\psi,\delta\psi))$
$=$ $d(\delta\mathbf{Q}_{\xi}-i_{\xi}\mathbf{\Theta}(\psi,\delta\psi))\,.$
To make contact with the first law of black hole thermodynamics, we take $\xi^{\mu}$ to be the timelike Killing vector that becomes null on the horizon. Wald shows that the variation of the Hamiltonian $\mathcal{H}$ with respect to the integration constants of a specific solution $\psi$ is given by
$\delta{\cal H}=\frac{1}{2\kappa}\int_{C}\mathbf{\omega}(\psi,\delta\psi,\mathscr{L}_{\xi}\psi)=\frac{1}{2\kappa}\int_{\Omega_{D-2}}(\delta\mathbf{Q}_{\xi}-i_{\xi}\mathbf{\Theta}(\psi,\delta\psi))\,$, (64)
where $C$ denotes a Cauchy surface, and $\Omega_{D-2}$ is its boundary, which has two components, one at infinity and one on the horizon. Thus according to the Wald formalism, the first law of black hole thermodynamics is a consequence of
$\delta{\cal H}_{\infty}=\delta{\cal H}_{r_{h}}\,.$ (65)
For $D=4$ dimensional EbM theory with $L=L_{1}+L_{2}$, we have
$\Theta^{\mu}=\Theta^{\mu}_{1}+\Theta^{\mu}_{2}\,$, (66)
with
$\Theta^{\mu}_{1}$ $=$ $g^{\mu\rho}g^{\nu\sigma}(\nabla_{\sigma}\delta g_{\nu\rho}-\nabla_{\rho}\delta g_{\nu\sigma})+\frac{\gamma}{2}\Big[2g^{\mu\lambda}B^{\rho}B^{\sigma}\nabla_{\rho}\delta g_{\lambda\sigma}$ (67)
$-g^{\mu\lambda}B^{\rho}B^{\sigma}\nabla_{\lambda}\delta g_{\rho\sigma}-g^{\rho\lambda}B^{\mu}B^{\sigma}\nabla_{\sigma}\delta g_{\rho\lambda}-2g^{\rho\lambda}\nabla_{\rho}(B^{\mu}B^{\sigma})\delta g_{\sigma\lambda}$
$+g^{\mu\lambda}\nabla_{\lambda}(B^{\rho}B^{\sigma})\delta g_{\rho\sigma}+g^{\rho\lambda}\nabla_{\sigma}(B^{\mu}B^{\sigma})\delta g_{\rho\lambda}\Big]-2\kappa B^{\mu\nu}\delta B_{\nu}\,$,
$\Theta^{\mu}_{2}$ $=$ $-\left[(1-4\gamma_{1}B^{\lambda}B_{\lambda})F^{\mu\nu}+4\gamma_{2}F^{\lambda[\mu}B^{\nu]}B_{\lambda}\right]\delta A_{\nu}=-{\cal F}^{\mu\nu}\delta A_{\nu}\,.$ (68)
Specializing to a variation induced by $\xi^{\mu}$, and substituting the following equations
$\delta_{\xi}g_{\mu\nu}$ $=$ $\nabla_{\mu}\xi_{\nu}+\nabla_{\nu}\xi_{\mu}\,$, (69)
$\delta_{\xi}B_{\mu}$ $=$ $\xi^{\rho}\nabla_{\rho}B_{\mu}+B_{\rho}\nabla_{\mu}\xi^{\rho}\,$, (70)
$\delta_{\xi}A_{\mu}$ $=$ $\xi^{\rho}\nabla_{\rho}A_{\mu}+A_{\rho}\nabla_{\mu}\xi^{\rho}\,$, (71)
into Eq. (66), the Noether charge $Q_{\xi}^{\mu\nu}=Q_{1}^{\mu\nu}+Q_{2}^{\mu\nu}$ can be obtained from
$\nabla_{\nu}Q_{\xi}^{\mu\nu}=J_{\xi}^{\mu}=\Theta^{\mu}(\delta_{\xi})-\xi^{\mu}L+\textup{EOM}\,$, (72)
and is given by
$Q_{1}^{\mu\nu}$ $=$ $-2\nabla^{[\mu}\xi^{\nu]}+2\gamma\left[\xi_{\lambda}\nabla^{[\mu}(B^{\nu]}B^{\lambda})-\xi^{[\mu}\nabla_{\lambda}(B^{\nu]}B^{\lambda})-B^{\lambda}B^{[\mu}\nabla_{\lambda}\xi^{\nu]}\right]-2\kappa B^{\mu\nu}B^{\lambda}\xi_{\lambda}\,$, (73)
$Q_{2}^{\mu\nu}$ $=$ $-{\cal F}^{\mu\nu}A^{\lambda}\xi_{\lambda}\,.$ (74)
To specialize our black hole ansatz (13)–(16), the result for the gravity part is well established and is given by
$\mathbf{Q}_{\xi 1}$ $=$ $r^{2}\left[\frac{h^{\prime}\sqrt{f}}{\sqrt{h}}+\frac{\ell\sqrt{f}\left(rh^{\prime}-4h\right)}{2r\sqrt{h}}\right]du\wedge d\varphi\,$, (75)
$i_{\xi}\mathbf{\Theta}_{1}$ $=$ $r^{2}\Bigg[\left(\frac{2\delta f\sqrt{h}}{r\sqrt{f}}+\frac{\delta h^{\prime}\sqrt{f}}{\sqrt{h}}-\frac{h^{\prime}\sqrt{f}\delta h}{2h^{3/2}}+\frac{h^{\prime}\delta f}{2\sqrt{fh}}\right)$ (76)
$+\ell\left(\frac{\delta f\sqrt{h}}{r\sqrt{f}}-\frac{\delta h\sqrt{f}}{r\sqrt{h}}+\frac{\delta fh^{\prime}}{4\sqrt{hf}}+\frac{\sqrt{f}\delta h^{\prime}}{2\sqrt{h}}-\frac{\delta h\sqrt{f}h^{\prime}}{4h^{3/2}}\right)\Bigg]du\wedge d\varphi\,$,
$\delta\mathbf{Q}_{\xi 1}-i_{\xi}\mathbf{\Theta}_{1}$ $=$ $r^{2}\Bigg[-\frac{2\delta f\sqrt{h}}{r\sqrt{f}}(1+\ell)\Bigg]du\wedge d\varphi\,.$ (77)
These results are consistent with previous works [49, 19]. For the electromagnetic part, it is convenient to introduce a total derivative term $d(\Psi\delta A_{{{\scriptscriptstyle(1)}}})$ [99, 104, 119] to the integrand of $\delta{\cal H}$. This term does not affect the final result, since $\delta\mathbf{Q}_{\xi}-i_{\xi}\mathbf{\Theta}$ is closed, and is introduced to circumvent the Dirac string singularity that would otherwise appear in $\delta\mathcal{H}$ due to the presence of a magnetic charge. The scalar $\Psi$ is defined via
$d\Psi=i_{\xi}*\mathcal{F}_{{{\scriptscriptstyle(2)}}}\quad\Rightarrow\quad\Psi=-i_{\xi}\mathcal{A}_{{{\scriptscriptstyle(1)}}}=-\frac{2p(\ell+1)^{3/2}}{(3\ell+2)r}\,.$ (78)
Thus we have
$\delta\mathbf{Q}_{\xi 2}=\delta(-*\mathcal{F}_{{{\scriptscriptstyle(2)}}}i_{\xi}A_{{{\scriptscriptstyle(1)}}})=-\delta(*\mathcal{F}_{{{\scriptscriptstyle(2)}}})i_{\xi}A_{{{\scriptscriptstyle(1)}}}-*\mathcal{F}_{{{\scriptscriptstyle(2)}}}\delta(i_{\xi}A_{{{\scriptscriptstyle(1)}}})\,$, (79)
$i_{\xi}\mathbf{\Theta}_{2}=i_{\xi}(-*\mathcal{F}_{{{\scriptscriptstyle(2)}}}\wedge\delta A_{{{\scriptscriptstyle(1)}}})=-i_{\xi}(*\mathcal{F}_{{{\scriptscriptstyle(2)}}})\wedge\delta A_{{{\scriptscriptstyle(1)}}}-*\mathcal{F}_{{{\scriptscriptstyle(2)}}}\delta(i_{\xi}A_{{{\scriptscriptstyle(1)}}})\,$, (80)
$d(\Psi\delta A_{{{\scriptscriptstyle(1)}}})=d\Psi\delta A_{{{\scriptscriptstyle(1)}}}+\Psi\delta dA_{{{\scriptscriptstyle(1)}}}=i_{\xi}(*\mathcal{F}_{{{\scriptscriptstyle(2)}}})\wedge\delta A_{{{\scriptscriptstyle(1)}}}-i_{\xi}\mathcal{A}_{{{\scriptscriptstyle(1)}}}\delta F_{{{\scriptscriptstyle(2)}}}\,$, (81)
$\delta\mathbf{Q}_{\xi 2}-i_{\xi}\mathbf{\Theta}_{2}-d(\Psi\delta A_{{{\scriptscriptstyle(1)}}})=-i_{\xi}A_{{{\scriptscriptstyle(1)}}}\delta(*\mathcal{F}_{{{\scriptscriptstyle(2)}}})+i_{\xi}\mathcal{A}_{{{\scriptscriptstyle(1)}}}\delta F_{{{\scriptscriptstyle(2)}}}\,.$ (82)
Finally, we have
$\delta{\cal H}=\frac{w_{2}}{\kappa}\sqrt{1+\ell}\delta m\,$, (83)
At infinity, we have
$\delta{\cal H}_{\infty}=\delta M=\frac{w_{2}}{\kappa}\sqrt{1+\ell}\delta m\,$, (84)
which indicates that the mass $M$ is
$M=\frac{mw_{2}}{\kappa}\sqrt{1+\ell}\,$, (85)
At horizon, we have
$\delta{\cal H}_{r_{h}}=T\delta S+\Phi_{e}\delta Q_{e}+\Phi_{m}\delta Q_{m}\,$, (86)
and the entropy $S$ can be easily calculated by
$S=\frac{2\pi w_{2}r_{h}^{2}(1+\ell)}{\kappa}\,.$ (87)
Thus the first law of black hole thermodynamics reads
$\delta M=T\delta S+\Phi_{e}\delta Q_{e}+\Phi_{m}\delta Q_{m}\,.$ (88)
The integral first law of black hole thermodynamics, also called Smarr formula, is given by
$M=2TS+\Phi_{e}Q_{e}+\Phi_{m}Q_{m}\,.$ (89)
When the topological parameter $k=0$, corresponding to planar black holes, there exists an additional generalized Smarr relation [120]
$M=\frac{2}{3}(TS+\Phi_{e}Q_{e}+\Phi_{m}Q_{m})\,.$ (90)
It is evident that, although the Lorentz-violating parameter $\ell$ affects nearly all thermodynamic quantities of the dyonic RN-like black hole—except for the electric potential and magnetic charge—all forms of the first law of black hole thermodynamics, Eqs. (88)–(90), remain consistent with their GR counterparts.
V Dyonic Taub-NUT-like black holes
Owing to the Misner string singularity, the Taub–NUT solution [121, 122] constitutes a nontrivial generalization of conventional static black holes.
V.1 Exact solution
The bumblebee theory we consider (27) admits a Taub-NUT dyonic solution
$ds^{2}$ $=$ $-h(r)(dt+2Nud\varphi)^{2}+\frac{dr^{2}}{f(r)}+(r^{2}+N^{2})\Big(\frac{du^{2}}{1-ku^{2}}+(1-ku^{2})d\varphi^{2}\Big)\,$, (91)
$A_{{{\scriptscriptstyle(1)}}}$ $=$ $\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[\phi(r)(dt+2Nud\varphi)+pud\varphi\Big]+c_{e}dt\,,\qquad B_{{{\scriptscriptstyle(1)}}}=b_{r}(r)dr\,$ (92)
where
$\phi(r)$ $=$ $\frac{1}{2N}\Bigg\{q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]-p\Bigg\}\,$, (93)
$b_{r}(r)$ $=$ $\frac{b}{\sqrt{f(r)}}\,,\qquad f(r)=\frac{h(r)}{1+\ell}\,,\qquad h(r)=\frac{2k(r^{2}-N^{2})-2mr+\frac{p^{2}+q^{2}}{\ell+2}}{2(r^{2}+N^{2})}\,,\qquad\ell=b^{2}\gamma\,.$ (94)
For $N=0$, the solution reduces to that presented in Section III. For convenience, we redefine the magnetic charge parameter $p$, which differs from the one in (14) by a constant factor $\sqrt{(3\ell+2)/((\ell+1)(\ell+2))}$. The parameter $c_{e}$ is a gauge constant, which does not affect the solution. In previous RN-like solution, it was set to zero.
Unlike conventional black holes, the Taub-NUT solution involves a Misner string singularity, and its thermodynamic analysis remains debated. In this work, we adopt the approach of [105] to evaluate the thermodynamics. We begin with the thermodynamic quantities that are well-defined, and then proceed to evaluate those requiring special treatment. In order to compute a well-define NUT potential, we can only handle the sphere topology $k=1$ case.
V.2 Normal thermodynamic quantities
The temperature $T$ of the Taub–NUT black hole is still determined by the surface gravity $K$, as expressed in (36)
$T=\frac{1}{4\pi\sqrt{\ell+1}}\frac{1}{r_{h}}\Big[1-\frac{p^{2}+q^{2}}{2(\ell+2)(r_{h}^{2}+N^{2})}\Big]\,.$ (95)
Using the expression for the Maxwell field $A_{{{\scriptscriptstyle(1)}}}$ and its EOM, one can determine the electric potential $\Phi_{e}$ and the electric charge $Q_{e}$. Likewise, the magnetic charge $Q_{m}$ (43) can be obtained via the Bianchi identity $dF_{{{\scriptscriptstyle(2)}}}=0$, and the magnetic potential $\Phi_{m}$ (45) can be calculated by introducing the dual 1-form field $\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ (44)
$\mathcal{A}_{{{\scriptscriptstyle(1)}}}$ $=$ $\mathcal{A}_{t}dt+\mathcal{A}_{\varphi}d\varphi+c_{m}dt\,$, (96)
$\mathcal{A}_{t}$ $=$ $-\frac{\sqrt{\ell+1}}{2\kappa N(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (98)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}\,$,
$\mathcal{A}_{\varphi}$ $=$ $-\frac{u\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (100)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}\,.$
Here, $c_{m}$ is a gauge parameter, which, similarly to the RN-like case, we also set to zero. The electric and magnetic potentials (42,45) are given by
$\Phi_{e}$ $=$ $\frac{1}{2N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (102)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$\Phi_{m}$ $=$ $\frac{\sqrt{\ell+1}}{2\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (104)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
Because of the Misner string, computing the electric charge requires careful treatment of the integration. Here,
$Q_{e}$ $=$ $\frac{1}{2\kappa}\int\mathcal{Q}_{tr}(r)dt\wedge dr+\mathcal{Q}_{r\varphi}(r,u)dr\wedge d\varphi+\mathcal{Q}_{u\varphi}(r)du\wedge d\varphi\,$, (105)
$\mathcal{Q}_{r\varphi}$ $=$ $-2Nu\mathcal{Q}_{tr}\,$, (106)
$\mathcal{Q}_{tr}$ $=$ $\frac{\ell+1}{\kappa\sqrt{(\ell+2)(3\ell+2)}(r^{2}+N^{2})}\Bigg\{p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (108)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$\mathcal{Q}_{u\varphi}$ $=$ $-\frac{\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (110)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
Following the integration procedure outlined in [105], the electric charge $Q_{e}$
$Q_{e}$ $=$ $\frac{1}{2\kappa}\int d\varphi\Big[\int_{-1}^{1}\mathcal{Q}_{u\varphi}(r\to\infty)du^{\prime}+\int_{r_{h}}^{r\to\infty}\mathcal{Q}_{r\varphi}(r^{\prime},u)dr^{\prime}\Big]=-\frac{w_{2}}{2}\mathcal{A}_{\varphi}|_{u=-1}^{u=1}$ (111)
$=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}\Bigg\{q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (113)
$-p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
can be determined. The magnetic charge is calculated using the same integration procedure
$Q_{m}$ $=$ $\frac{1}{2\kappa}\int\mathcal{P}_{tr}(r)dt\wedge dr+\mathcal{P}_{r\varphi}(r,u)dr\wedge d\varphi+\mathcal{P}_{u\varphi}(r)du\wedge d\varphi\,$, (114)
$\mathcal{P}_{r\varphi}$ $=$ $-2Nu\mathcal{P}_{tr}\,$, (115)
$\mathcal{P}_{tr}$ $=$ $-\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\phi^{\prime}(r)\,,\qquad\mathcal{P}_{u\varphi}=\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\big(p+2N\phi(r)\big)\,$, (116)
resulting in $Q_{m}$ given by
$Q_{m}$ $=$ $\frac{1}{2\kappa}\int d\varphi\Big[\int_{-1}^{1}\mathcal{P}_{u\varphi}(r\to\infty)du^{\prime}+\int_{r_{h}}^{r\to\infty}\mathcal{P}_{r\varphi}(r^{\prime},u)dr^{\prime}\Big]=\frac{w_{2}}{2}A_{\varphi}|_{u=-1}^{u=1}$ (117)
$=$ $w_{2}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (119)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
V.3 Unnormal thermodynamic quantities
Although the calculations are involved, the evaluation of the electric and magnetic potentials and charges follows standard procedures. In contrast, the computation of the mass and NUT charge remains debated. In this work, we employ the method of [105] to provide a set of Taub–NUT black hole thermodynamic quantities that satisfy the first law.
To simplify the analysis, we introduce a scalar $\Psi$, following (78),
$\Psi$ $=$ $-\frac{\sqrt{\ell+1}}{\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (121)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}+c_{m}\,$,
to eliminate the Dirac string singularity associated with the magnetic charge. Owing to the presence of the Misner string, the Wald formalism diverges at $u=\pm 1$ during integration. We adopt the integration path illustrated in Fig. 1. Equation (64) can be separated into four distinct parts
$\delta\mathcal{H}=0=\delta\mathcal{H}_{S_{2}}+\delta\mathcal{H}_{S_{1}}+\delta\mathcal{H}_{T_{N}}+\delta\mathcal{H}_{T_{S}}\,.$ (122)
Figure 1: Owing to the NUT charge, the Wald formalism has singularities at $u=\pm 1$. We select the integration path illustrated below to bypass these singularities.
As discussed in Sec. IV, the Wald entropy formula fails to yield a black hole entropy consistent with the first law of thermodynamics $r=r_{h}$. Consequently, we set $S_{2}$ at the black hole horizon and, employing the Wald formalism [95], extract a self-consistent expression for the entropy from $\delta\mathcal{H}_{S_{1}}$. We choose the gauge parameters
$c_{e}=-\Phi_{e}\,,\qquad c_{m}=2\Phi_{m}\,$, (123)
such that $\Phi_{e}\delta Q_{e}+\Phi_{m}\delta Q_{m}$ appears at infinity, in which case $\delta\mathcal{H}_{r_{h}}$ yields the black hole entropy
$\delta\mathcal{H}_{r_{h}}=T\delta S\,,\qquad\Rightarrow\qquad S=\frac{2\pi w_{2}(\ell+1)}{\kappa}(r_{h}^{2}+N^{2})\,.$ (124)
For the Taub-NUT solution, besides the event horizon at $r=r_{h}$, there are also two Killing horizons located at the north and south poles $u=\pm 1$. The two degenerate Killing vectors are
$u=\pm 1:\qquad l_{\pm}=\partial_{\varphi}\mp 4\Phi_{N}\partial_{t}\,.$ (125)
At $u=\pm 1$, $l_{\pm}$ becomes a null vector $l_{\pm}^{2}=0$, which defines the NUT potential $\Phi_{N}$
$\Phi_{N}=\frac{N}{2}\,.$ (126)
For k=0,1k=0,-1 case, there is no north or south pole, which means that there is no straightforward way to define the NUT potential by the degenerate Killing vector.
The NUT parameter can give rise to corresponding electric and magnetic charges
$Q_{eN}$ $=$ $\int_{r=r_{h}}^{r\to\infty}\mathcal{F}_{tr}dr=2w_{2}\mathcal{A}_{t}|_{r\to\infty}^{r=r_{h}}$ (127)
$=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (129)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$Q_{mN}$ $=$ $\int_{r=r_{h}}^{r\to\infty}F_{tr}dr=2w_{2}A_{t}|_{r\to\infty}^{r=r_{h}}$ (130)
$=$ $\frac{w_{2}}{N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (132)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
These charges are also associated with corresponding conjugate thermodynamic potentials
$\Phi_{eN}$ $=$ $\frac{1}{8}l^{\mu}_{-}\Big(A_{\mu}(u=1)+A_{\mu}(u=-1)\Big)\bigg|^{r=r_{h}}_{r\to\infty}$ (133)
$=$ $\frac{1}{4}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (135)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$\Phi_{mN}$ $=$ $-\frac{1}{8}l^{\mu}_{-}\Big(\mathcal{A}_{\mu}(u=1)+\mathcal{A}_{\mu}(u=-1)\Big)\bigg|^{r=r_{h}}_{r\to\infty}$ (136)
$=$ $-\frac{\sqrt{\ell+1}}{4\kappa(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (138)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
We find that there is a relationship between the electric and magnetic charges (113, 119) and the charges induced by the NUT parameter
$Q_{e}=\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}q-2\Phi_{N}Q_{eN}\,,\qquad Q_{m}=w_{2}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}p+2\Phi_{N}Q_{mN}\,.$ (139)
If we take $S_{2}$ to be at infinity, then according to [95], we can set $c_{e}=0=c_{m}$ so that the electric and magnetic potentials appear on the horizon, in which case $\delta\mathcal{H}_{S_{2}}$ gives the black hole mass
$-\delta\mathcal{H}_{\infty}=\frac{w_{2}\sqrt{\ell+1}}{\kappa}\delta\mu\,.$ (140)
However, as noted in [105], if the above expression were taken as the mass, the black hole mass could be either positive or negative. Since a non-positive-definite mass has no physical meaning, the expression represents only a partial contribution to the total mass. We find that the charges (132) and the potentials (138) induced by the NUT parameter can be consistently incorporated into the Wald formalism
$-\delta\mathcal{H}_{\infty}-\mathcal{H}_{T_{N}}-\delta\mathcal{H}_{T_{S}}=\delta M-\Phi_{N}\delta Q_{N}-\Phi_{eN}\delta Q_{eN}-\Phi_{mN}\delta Q_{mN}\,.$ (141)
This procedure yields well-defined expressions for the mass $M$ and the NUT charge $Q_{N}$
$M$ $=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa}\mu+2\Phi_{N}Q_{N}\,$, (142)
$Q_{N}$ $=$ $\frac{2w_{2}N\sqrt{\ell+1}}{\kappa r_{h}}\Bigg\{1-\frac{p^{2}+q^{2}}{4N^{3}(\ell+2)}\Big(2N$ (144)
$-r_{h}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Big)\Bigg\}\,.$
The Wald formalism (122) provides the first law of thermodynamics for the Taub–NUT solution
$\delta M=T\delta S+\Phi_{e}\delta Q_{e}+\Phi_{m}\delta Q_{m}+\Phi_{N}\delta Q_{N}+\Phi_{eN}\delta Q_{eN}+\Phi_{mN}\delta Q_{mN}\,.$ (146)
The Smarr relation is
$M=2TS+\Phi_{e}Q_{e}+\Phi_{m}Q_{m}\,.$ (147)
At this stage, we have completed the construction of the dyonic Taub-NUT-like black hole solution and the derivation of its first law of thermodynamics. On one hand, the exact dyonic Taub-NUT-like solution further demonstrates the internal consistency and robustness of the specific EbM model. On the other hand, although the presence of the Taub-NUT charge introduces nontrivial effects on the black hole thermodynamics, the quantities computed via the Wald formalism still guarantee that the first law of black hole thermodynamics holds.
VI Generalization to higher dimensions
In previous sections, we obtained the dyonic RN-like black hole solutions in the four-dimensional EbM theory, and derived the first law of thermodynamics for these black holes. Now in this section, we will generalize these results to arbitrary even dimensions $D=2+2n$.
VI.1 High dimensional dyonic RN-like solutions
The general ansatz for dyonic topological black holes in $D=2+2n$ dimensions is given by
$ds^{2}$ $=$ $-h(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}\sum_{i=1}^{n}d\Omega_{i,k}^{2}\,$, (148)
$F$ $=$ $-\phi^{\prime}(r)dt\wedge dr+p\sum_{i=1}^{n}dx_{i}\wedge dy_{i}\,$, (149)
$B$ $=$ $b_{r}(r)dr\,$, (150)
with
$d\Omega_{i,k}^{2}=\frac{dx_{i}^{2}}{1-kx_{i}^{2}}+(1-kx_{i}^{2})dy_{i}^{2}\,.$ (151)
The solutions for are
$h$ $=$ $k-\frac{m}{r^{2n-1}}+\frac{(2n-1)q^{2}}{2\left(2n+\ell\right)r^{2(2n-1)}}+\frac{n(2n-1)\left(1+\ell\right)p^{2}}{2(3-2n)\left(2n+(2n+1)\ell\right)r^{2}}\,$,
$f$ $=$ $\frac{h}{(2n-1)(1+\ell)}\,,\quad\phi=\frac{q}{r^{2n-1}}\,,\quad b_{r}=\frac{b}{\sqrt{f}}\,$, (152)
where
$\gamma_{1}=\frac{\gamma}{4(2n+(2n+1)\ell)}\,,\quad\gamma_{2}=-\frac{2n^{2}(1+\ell)\gamma}{(2n+\ell)(2n+(2n+1)\ell)}\,.$ (153)
The horizon topology now becomes ${\cal M}_{2}\times{\cal M}_{2}\times\dots\times{\cal M}_{2}$, where ${\cal M}_{2}$ can be sphere, torus, or hyperbolic 2-space.
VI.2 Thermodynamics
Here we present the thermodynamical properties of the $D=2+2n$-dimensional dyonic RN-like topological black holes
$m=kr_{h}^{2n-1}-\frac{n(2n-1)(\ell+1)p^{2}r_{h}^{2n-3}}{2(2n-3)((2n+1)\ell+2n)}+\frac{(2n-1)q^{2}}{2(\ell+2n)r_{h}^{2n-1}}\,.$
$T=\frac{h^{\prime}}{4\pi\sqrt{(2n-1)(1+\ell)}}=\frac{k\sqrt{2n-1}}{4\pi r_{h}\sqrt{\ell+1}}-\frac{n\sqrt{2n-1}p^{2}\sqrt{\ell+1}}{8\pi r_{h}^{3}(2n+(2n+1)\ell)}-\frac{(2n-1)^{3/2}q^{2}}{8\pi\sqrt{\ell+1}(2n+\ell)r_{h}^{4n-1}}\,$,
$Q_{e}=\frac{n\sqrt{(2n-1)(\ell+1)}qw_{2}^{n}}{(\ell+2n)\kappa}\,,\quad\Phi_{e}=\frac{q}{r_{h}^{2n-1}}\,$,
$Q_{m}=\frac{npw_{2}}{2\kappa}\,,\quad\Phi_{m}=-\frac{2n\sqrt{2n-1}p(\ell+1)^{3/2}w_{2}^{n-1}r_{h}^{2n-3}}{(2n-3)((2n+1)\ell+2n)}\,$,
$M=\frac{nmw_{2}^{n}}{\sqrt{2n-1}\kappa}\sqrt{\ell+1}\,,\quad S=\frac{2\pi r_{h}^{2n}(\ell+1)w_{2}^{n}}{\kappa}\,.$ (154)
The differential version of first law of black hole thermodynamics can also be written as Eq. (88), while the integral form reads
$M=\frac{2n}{2n-1}TS+\Phi_{e}Q_{e}+\frac{1}{2n-1}\Phi_{m}Q_{m}\,.$ (155)
For the case with $k=0$, the generalized Smarr relation takes the form
$M=\frac{2n}{2n+1}(TS+\Phi_{e}Q_{e})+\frac{2}{2n+1}\Phi_{m}Q_{m}\,.$ (156)
All forms of the first law of black hole thermodynamics in higher dimensions remain consistent with those in GR.
VI.3 Some explicit examples
In the purely electric case, i.e., $p=0$, we can first set $k=0$ and then rewrite $\sum_{i=1}^{n}d\Omega_{i,k}^{2}$ as $d\tilde{\Omega}_{D-2,k}^{2}$, where k=1,0,1k=1,0,-1 correspond to (D2)(D-2)-dimensional spherical, planar, and hyperbolic geometries, respectively. The arbitray $D$-dimensional electrically charged RN-like topological black hole solution is given by
$h$ $=$ $k-\frac{m}{r^{D-3}}+\frac{(D-3)q^{2}}{2\left((D-2)+\ell\right)r^{2(D-3)}}\,$,
$f$ $=$ $\frac{h}{(D-3)(1+\ell)}\,,\quad\phi=\frac{q}{r^{D-3}}\,,\quad b_{r}=\frac{b}{\sqrt{f}}\,$, (157)
where
$\gamma_{1}=\frac{\gamma}{4((D-2)+(D-1)\ell)}\,,\quad\gamma_{2}=-\frac{(D-2)^{2}(1+\ell)\gamma}{2((D-2)+\ell)((D-2)+(D-1)\ell)}\,.$ (158)
In the purely magnetic case, i.e., $q=0$, the only changes in Eq. (152) are
$h=k-\frac{m}{r^{2n-1}}+\frac{n(2n-1)\left(1+\ell\right)p^{2}}{2(3-2n)\left(2n+(2n+1)\ell\right)r^{2}}\,,\quad\phi=0\,.$ (159)
VII Conclusion
In this paper, we construct four-dimensional dyonic RN-like black holes with general topological horizons within Einstein-bumblebee gravity, one of the simplest vector-tensor theories realizing spontaneous Lorentz symmetry breaking. We then investigate their thermodynamic properties and employ the Wald formalism to compute the conserved mass and entropy, thereby establishing the first law of black hole thermodynamics. Furthermore, we extend the static dyonic solution to the Taub–NUT case, revealing its nontrivial black hole thermodynamics. Finally, we generalize the four-dimensional dyonic RN-like results to higher dimensions. These results provide a concrete framework for further studies of the effects of spontaneous Lorentz symmetry breaking in black hole physics and astrophysics.
Acknowledgements.
We are grateful to Yu-Qi Chen, Xing-Hui Feng, Zhi-Chao Li, Hai-Shan Liu, Yuxiao Liu, Yanzhao Lu, H. Lü and Hongwei Yu for useful discussions. S.L., and L.L. were supported in part by the National Natural Science Foundation of China (No. 12105098, No. 12481540179, No. 12075084, No. 11690034, No. 11947216, and No. 12005059) and the Natural Science Foundation of Hunan Province (No. 2022JJ40264), and the innovative research group of Hunan Province under Grant No. 2024JJ1006. L.M. is also supported in part by National Natural Science Foundation of China (NSFC) grant No. 12447138, Postdoctoral Fellowship Program of CPSF Grant No. GZC20241211 and the China Postdoctoral Science Foundation under Grant No. 2024M762338.